# Zhiwei Zuo Zhuo Tang Kenli Li Anwitaman Datta

### v2

@@ 4.3.3 Effectiveness of Top-K and Random-k
4.3.4 Mixed Top-K and Random-k
In Random-k, parameters are chosen randomly for perturbation, while in Top-K, parameters to be perturbed are carefully calculated. To investigate whether the random selection in Random-k is efficient enough or can be further improved, we designed a set of experiments in which $K$ in $k\%$ parameter selected by Random-k is replaced by those chosen by Top-K. This formed a mixed Random-TopK unlearning strategy, and related experiments are conducted using settings similar to the aforementioned ones. The $Acc_{RE}$ for Top-K, Random-k, and Random-TopK after 100 training epochs are shown in Fig.8.
Based on the $Acc_{RE}$ trends for Top-K, Random-K, and Random-TopK, at the beginning of training, the $Acc_{RE}$ for Top-K is higher than that for Random-K and Random-TopK. This implies that as the number of perturbed parameters in the model increases, more knowledge is eliminated. For Random-TopK, its $Acc_{RE}$ is lower than both Top-K and Random-K, indicating that over-modifying parameter based on those with the highest sensitivity will be more time-consuming. For models like ResNet18 (Fig.8(a)-8(d)), VGG (Fig.8(e)-8(h)), and DenseNet (Fig.8(i)-8(l)), Random-TopK still exhibits a certain upward trend in $Acc_{RE}$, suggesting that even with excessive parameter modifications, Random-TopK can still achieve partial unlearning effects. However, we notice that when using the Random-Topk stragegy on GoogLeNet (Fig.8(m)-8(p)), the model fell into a state of complete non-learning. We believe this may be due to the different sensitivities and parameter dependencies between the different branches of the Inception module in GoogLeNet. Employing the Random-TopK strategy might introduce complexities, adversely affecting certain branches.
4.3.5 Unlearning Degree Evaluation
@@
The resulting unlearning degree is illustrated in Fig.9. The uniform random output accuracy for CIFAR-10 is around 10%, therefore, when we conduct experiments with CIFAR-10, the range of unlearning degree is [0, 90%]. From the experimental results in Fig.9, it can be observed that the use of Top-K achieves deepest unlearning degree that close to 90%. Random-k follows, achieving the second-deepest unlearning degree. However, due to EU-K and CF-K freezing most of parameters and modifying the last $K$ layers, their degree of unlearning is significantly lower than that of Top-K and Random-k. Moreover, at lower ratio of unlearning data, such as 5%, the degree of unlearning across different unlearning strategies if relatively higher.
5 Conclusion
In this paper, we propose fine-grained model weights perturbation methods to achieve inexact machine unlearning. Our method ensures a higher degree of unlearning the information from dataset $D_{UL}$ while maintaining model indistinguishability and achieving largest acceleration effect. Furthermore, after using SPD-GAN to break the i.i.d attribute of the unlearning data, we can quantify the degree of unlearning. The results further demonstrate that our proposed Top-K based inexact unlearning strategy achieves the best unlearning effect.
Limitations and Future Work
Although the Top-K strategy can achieve the best unlearning effect with the smallest perturbation ratio and the fastest acceleration, our current analysis of parameter sensitivity in the model ignores the dependence between parameters. One possible direction is to redesign and select Top-K parameters for unlearning process considering the dependencies between parameters. By examining these dependencies during the unlearning process, it would become possible to interpret and explain the effects of unlearning on different components of the model, enhancing interpretability and transparency. This, in turn, is expected to facilitate better optimization of the unlearning process to improve overall model performance.
@@ Acknowledgments
The work is supported by the National Key Research and Development Program of China (2021ZD40303), the National Natural Science Foundation of China (Grant Nos. 62225205, 92055213), Natural Science Foundation of Hunan Province of China (2021JJ10023), Shenzhen Basic Research Project (Natural Science Foundation) (JCYJ20210324140002006).

### v3

@@ IV-C3 Effectiveness of Top-K and Random-k
IV-C4 Mixed Top-K and Random-k
In Random-k, parameters are chosen randomly to be perturbed, while in Top-K, parameters to be perturbed are carefully calculated. To investigate whether the random selection in Random-k is efficient enough or can be further improved, we designed a set of experiments in which $K$ in $k\%$ parameters selected by Random-k is replaced by those chosen by Top-K. This formed a mixed Random-TopK unlearning strategy, and related experiments are conducted using settings similar to the aforementioned ones. The $Acc_{RE}$ for Top-K, Random-k, and Random-TopK after 100 training epochs are shown in Fig.8.
Based on the $Acc_{RE}$ trends for Top-K, Random-k, and Random-TopK, at the beginning of training, the $Acc_{RE}$ for Top-K is higher than that for Random-k and Random-TopK. This implies that as the number of perturbed parameters in the model increases, more knowledge is eliminated. For Random-TopK, its $Acc_{RE}$ is lower than both Top-K and Random-k, indicating that over-modifying parameter based on those with the highest sensitivity will be more time-consuming. For models like ResNet18 (Fig.8(a)-8(d)), VGG (Fig.8(e)-8(h)), and DenseNet (Fig.8(i)-8(l)), Random-TopK still exhibits a certain upward trend in $Acc_{RE}$, suggesting that even with excessive parameter modifications, Random-TopK can still achieve partial unlearning effects. However, we notice that when using the Random-Topk stragegy on GoogLeNet (Fig.8(m)-8(p)), Random-TopK still exhibits a certain upward trend in $Acc_{RE}$, suggesting that even with excessive parameter modifications, Random-TopK can still achieve partial unlearning effects. However, we notice that when using the Random-Topk stragegy on GoogLeNet (Fig.8(m)-8(p)), the model fell into a state of complete non-learning. We believe this may be due to the different sensitivities and parameter dependencies between the different branches of the Inception module in GoogLeNet. Employing the Random-TopK strategy might introduce complexities, adversely affecting certain branches.
IV-C5 Unlearning Degree Evaluation
@@
The resulting unlearning degree is illustrated in Fig.9. The uniform random output accuracy for CIFAR-10 is around 10%, therefore, when we conduct experiments with CIFAR-10, the range of unlearning degree is [0, 90%]. The degree of unlearning indicates, beyond the model generalization caused high $Acc_{UL}$, the extent to which the machine unlearning actually eliminates the influence of unlearning data on the model. From the experimental results in Fig.9, it can be observed that Top-K achieved the best degree of unlearning (peaking at 88.4% when 10% data was unlearned), followed closely by Random-k, which reached a suboptimal level of unlearning (peaking at 86% when 5% of the data was unlearned). However, due to EU-K and CF-K freezing most of parameters and modifying the last $K$ layers, their degree of unlearning is significantly lower than that of Top-K and Random-k. Moreover, at lower ratio of unlearning data, such as 5%, the degree of unlearning across different unlearning strategies if relatively higher. This implies that as more data needs to be forgotten, achieving better unlearning effects may become more difficult.
V Conclusion
In this paper, we propose fine-grained model weights perturbation methods to achieve inexact machine unlearning. Our method ensures a higher degree of unlearning the information from dataset $D_{UL}$ while maintaining model indistinguishability and achieving largest acceleration effect. Furthermore, we design a novel machine unlearning degree quantification method using SPD-GAN to break the i.i.d attribute of the unlearning data. The results further demonstrate that our proposed Top-K based inexact unlearning strategy achieves the best unlearning effect and reveal that it is more difficult to unlearn more data.
Limitations and Future Work Although the Top-K strategy can achieve the best unlearning effect with the smallest perturbation ratio, the fastest acceleration and the optimal unlearning degree, our current analysis of parameter sensitivity in the model ignores the dependence between parameters. One possible direction is to redesign and select Top-K parameters for unlearning process considering the dependencies between parameters. By examining these dependencies during the unlearning process, it would become possible to interpret and explain the effects of unlearning on different components of the model, enhancing interpretability and transparency. This, in turn, is expected to facilitate better optimization of the unlearning process to improve overall model performance.
@@ Acknowledgments
The work is supported by the National Natural Science Foundation of China (Grant Nos. 62225205, 92055213), the National Key Research and Development Program of China (2021ZD40303), the Science and Technology Program of Changsha (kh2301011), Shenzhen Basic Research Project (Natural Science Foundation) (JCYJ20210324140002006) and China Scholarship Council.